Chlordiazepoxide is used to ease the agitation associated with alcohol withdrawal. one form of asministration is by IM injection. the loading dose is 100 mg. the maintenance dose consists of 50 mg injections three times a day. the first days dose consists of the loading dose and two 50-mg maintenance injections. if the chlordiazepoxide solution has a concentration of 50 mg/ml, how many milliliters of solution will be needed for one week?
Answer #1

First day dosage

Loading dose=100 mg

Maintenance dose=50 mg * 2=100 mg

Total dosage=100+100=200 mg

Maintenance dose for 1 day= 50 mg * 3(times a day)

=150 mg/day

Maintenance dose for 6 days=150 mg * 6=900 mg

Total dosage in 1 week=200+900=1100 mg

Since 1 ml contains 50 mg,

1 ml=50 mg

X =1100 mg(X=unknown ml)

X *50=1100*1

X=1100/50=22 millilitres,So 22 ml of solution will be needed for 1 week.

On the first day a loading dose of 100 mg and a maintenance dose of 50 mg twice a day is given.So a total of 200 mg is given in the first day of the week.

For calculating the dose for the next 6 days,the dose for 1 day need to be find out which is 50 multiplied by 3(as it is given thrice a day) equals 150 mg per day.

So dose for 6 days is 150 mg multiplied 6 equals 900 mg

Total dose in 1 week is calculated as 200+900=1100 mg

Since 1 ml contains 50 mg,millilitres of solution needed for one week will be 1100 divided by 50 equals 22 millilitres

Therefore, 22 millilitres of solution will be needed for 1 week